A uniform magnetic field passes through a horizontal circular wire loop at an angle 60° from the normal to plane of the loop. The magnitude of the magnetic field is 3.0 T, and the radius of the wire loop is 10 cm. Find the magnetic flux through the loop.
Answer #1

magnetic flux = BA cos \theta

where B = magneitc field

A = Area , \theta= angle between magnetic field and area vector

= 3*pi*0.1^2*cos 60

= 0.047 T-m^2

so the magnetic flux through the loop is 0.047 T-m^2